A coffee shop wants to spice up their standard blend of coffee by mixing in cinnamon spiced coffee beans. Their standard blend costs $7.95 per pound and the cinnamon coffee costs $12.57 per pound. They want the mix to cost $10.33 per pound, and they need 47 pounds of the mix. How many pounds of cinnamon coffee beans do they need to make this mix?
Answer by mananth(16946)

price ---------------- quantity
Standard blend $7.95 ---------------- x pounds
cinnamon based $12.57 ------ 47 - x pounds
Mixture $10.33 ---------------- 47

7.95 x + 12.57 ( 47 - x ) = 485.51
7.95 x + 590.79 - 12.57 x = 485.51
7.95 x - 12.57 x = 485.51 - -590.79
-4.62 x = -105.28
/ -4.62
x = 22.79 pounds $7.95 Standard blend
24.21 pounds $10.33 cinnamon based

m.ananth@hotmail.ca